Mary receives $87.00 for working one eight our day. One day she had to stop after working five hours because of a doctor's appointment. How much did she make that day?
Answer by stanbon(75887)

Mary receives $87.00 for working one eight hour day. One day she had to stop after working five hours because of a doctor's appointment. How much did she make that day?
----------------
Using a proportion:
x/5 = 87/8
-----------
Cross-multiply to get:
x = 5(87/8)
---
x = 435/8
----
x = $54.38
=================================
Cheers,
Stan H.